OpenStudy (anonymous):

The square root of a number is 17. Which method can be used to find the value of this number

Parth (parthkohli):

The calculator method is what they call it.

Parth (parthkohli):

Just kidding, try to make up an equation and solve it.

OpenStudy (anonymous):

idk how to all i need is to know how to like Take the square of 17. or something so confused

Parth (parthkohli):

No, the square root of a number is \(17\).

Parth (parthkohli):

Now if you denote your number by \(x\), how would you denote "square root of \(x\)"?

OpenStudy (anonymous):

this is confusing i have no clue what your saying

Parth (parthkohli):

...

Parth (parthkohli):

Do you know what square root means?

OpenStudy (anonymous):

noppee

Parth (parthkohli):

...really?

OpenStudy (anonymous):

yesss :C

Parth (parthkohli):

The square root of a number \(n\) is \(k\) if \(k \times k = n\)

Parth (parthkohli):

So \(2\) is the square root of \(4\) because \(2 \times 2 = 4\). \(3\) is the square root of \(9\) because \(3 \times 3 = 9\).